OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95. 203 197 168 163

OpenStudy (anonymous):

you would have to know the formula used to find the nth term in an arithmetic sequence.. \[a_n=a_1+(n-1)d\]

OpenStudy (anonymous):

a(n)=the nth term a(1)=the first term of the sequence d=common difference

OpenStudy (campbell_st):

1st task is to find the common difference d, which can be done using the formula above and the 18th term 95 = -7 + ( 18 -1 )d which is 95 = -7 + 17d solve for d. then use the formula above with n = 35 hope this helps.